OpenStudy (anonymous):

The rate at which water flows out of a pipe in gallons per hour is given by a differentiable function R of time t. The table shows the rate as measured every 3 hours for a 24 hr period. Hours 0 R(t) 9.6 Hours 3 R(t) 10.4 Hours 6 R(t) 10.8 Hours 9 R(t) 11.2 Hours 12 R(t) 11.4 Hours 15 R(t) 11.3 Hours 21 R(t) 10.2 Hours 24 R(t) 9.6 The rate at which water flows out of a pipe, can be approximated by q(t)= 1/79(768+23t-t^2) Use q(t) to approximate the average rate of the water flow during the 24 hour time period. Indicate units of measure.

OpenStudy (anonymous):

using the equation q(t)=(1/79)[768+23t-t^2] we get the fol Hours 0 R(t)= 9.72 Hours 3 R(t) =10.48 Hours 6 R(t) =11.01 Hours 9 R(t) =11.32 Hours 12 R(t)= 11.39 Hours 15 R(t)= 11.24 Hours 21 R(t)= 10.25 Hours 24 R(t)= 9.42 get the total here to find the average ------------------- Rtotal=84.83 now average R=84.83/8=10.60 gal/hour ...ans

OpenStudy (anonymous):

use your calculator in the eq q(t)=(1/79)[768+23t-t^2]

OpenStudy (anonymous):

let me know if you have questions

OpenStudy (anonymous):

@mark_o thank you, but I don't know how to use my graphing calculator to figure it out.